Olga has 5 times as many dimes as nickels. If she has $3.30, how many of each coin does she have?
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

let n=nickels d=dimes
5n+10d=330
d=5n

Answer by stanbon(75887)

Olga has 5 times as many dimes as nickels. If she has $3.30, how many of each coin does she have?
-----------------
Equations:
d = 5n
5n + 10d = 330 cents
--------------------------
Substitute for d and solve for "n":
5n + 10(5n) = 330
5n + 50n = 330
55n = 330
n = 6 (# of nickels)
d = 5n = 5*6 30 (# of dimes)
================================
Cheers,
Stan H.